OpenStudy (anonymous):

Medal!! Write the equation of the transformed graph of tangent with period 4 that has been shifted vertically up 3 units.

OpenStudy (anonymous):

can you help @Loser66

OpenStudy (loser66):

I think it is \(y = 3+tan \dfrac{\pi}{4}x \)

OpenStudy (loser66):

since period of tangent is pi, your function has period 4, it means \(\dfrac{\pi}{pi/4}=4\) so, the "number" in the front of x must be \(\pi/4\). it is shifted 3 units up; it means the "whole" function + 3. So that, the equation is y = 3 + tan (pi/4) x

OpenStudy (anonymous):

oh ok I get that for the period. help on another? I'll make a new question so I can medal you